# Surrogate-based multilevel Monte Carlo methods for uncertainty quantification in the Grad-Shafranov free boundary problem

### v2

@@ 4 Numerical experiments
We used the discrete solution corresponding to these baseline currents as the initial value for Newton’s method when solving the perturbed problem. For every realization of $\boldsymbol{\omega}$, we made use of the finite element-based solver FEEQS.m [15, 22, 23], developed by H. Heumann and collaborators to obtain a numerical solution of (2).
@@ 4.1 Experiment description
Table 1: Number of spatial grid points $M_{\ell}$ at increasing spatial grid level $\ell=0$ to 5, and number of sparse grid points $P_{q}$ from sparse grid level $q=0$ to 4.
@@
For direct non-linear solves at each parameter value, the stopping threshold for the relative residual of Newton’s method is set to $5\times 10^{-11}$. The normalization factor $\|\mathbb{E}[u_{\ell}]\|_{Z}$ on the finest spatial mesh level ($\ell=5$) is approximately $8.5708\times 10^{-1}$. All experiments are conducted using MATLAB R2023a on a System 76 Thelio Major with 256GB RAM and a 64-Core @4.6 GHz AMD Threadripper 3 processor.
@@ 4.2 Surrogate construction (offline) costs
The cost of constructing the surrogate is then the cost of computing the direct solution on each sparse grid node times the number of sparse grid nodes. For these tests, the latter number is $25$, corresponding to sparse grid level $q=1$. We explored the cost of a single direct solution in [14]. These are summarized in the blue curve shown in the left image of Figure 2.(This part of the image is reproduced from [14].) They were obtained by solving the systems for 100 random currents and taking the mean CPU time, and they indicate that the costs were of the form $M^{\gamma}$ for $\gamma=1.09$. The resulting total costs are summarized in Table 3 and plotted in the right image of Figure 2, for both single-level and multilevel (in space) surrogates. The plots indicate that the two surrogates have asymptotic complexity of magnitude $\epsilon^{-1}$, which aligns with the theoretical cost $\epsilon^{-\gamma/\alpha}\approx\epsilon^{-1.09}$. The slightly larger cost of the multilevel version is due to the fact that it requires the surrogates for all coarser grids, as observed in the comments that follow (22).
@@
Figure 2: Left: Mean CPU times to compute 100 realizations of solutions for both direct and surrogate computations vs. an increasing number of spatial grid points $M_{\ell}$. Right: Offline costs of construction of both single spatial level and multi spatial level surrogates.
@@ 4.3 Sampling (online) costs
The MLMC-FE estimator for (3) involves accumulating surrogate sample corrections $\widehat{Y}_{\ell}^{(i)}$ across $L$ spatial levels. As we discussed in previous work [14], for each sample correction, $\widehat{u}_{\ell-1}^{(i)}$ is constructed on the coarse spatial grid by interpolating $\widehat{u}_{\ell}^{(i)}$ from the fine grid. This interpolation method serves as a more efficient, though slightly less accurate, alternative to the Galerkin projection. However, the use of non-nested geometry-conforming uniform spatial meshes introduces extrapolation errors during interpolation. This issue can be addressed using a sufficiently fine common grid (e.g., $\ell=5$) that encompasses all coarser meshes, in order to minimize or eliminate extrapolation errors caused by interpolation of surrogate corrections from all multi-level spatial coarser grids. Although this does increase accuracy, it does so at a significant extra cost. Table 4 provides a quantitative summary of the online CPU time for the various solution strategies, including interpolation to a common fine grid just discussed. The entries in this table correspond to the time required to make the square root of the sum of the squares of the discretization and statistical errors (see (24)) less than the tolerances, since the interpolation error is limited by the fixed choice of a sparse grid. The table uses the Monte Carlo method with direct computation as a benchmark and calculates the speedups of the different sampling strategies. It is evident that simply replacing the direct solve with the surrogate (third column) produces significant speedups (on the order of 10 to 50); replacing the full grids with multilevel grids (fourth column) results in somewhat better speedups (up to 200 for small tolerance), and using both surrogates and multilevel grids (fifth column) yields dramatic speedups, often exceeding $10^{4}$. The last two columns of the table present results for mitigating extrapolation errors using a common fine grid for MLMC with both direct solve and surrogate, revealing that the speedups for both approaches are comparable due to the significant interpolation cost. In the following, we discuss some issues regarding accuracy.
@@
where the term “$25C_{d}$” reflects the overhead of the offline costs for the level-1 collocation sparse grid. This means that the surrogate is less expensive if $k>(21\cdot 25)/20=26.25$. It is clear from Table 5 that typically many more samples are needed in practice. This comparison will also be affected by many factors: the direct solve entails solution of non-linear systems, which will be affected by the number of iterations, and as shown in Table 5, the grid sizes and number of samples will not in general be identical for the different strategies. Nevertheless, this simplified analysis clearly indicates the advantages of the surrogate for a simulation, despite its offline overhead.
@@ 4.4 Properties of geometric parameters
[…]

### v3

@@ 4 Numerical experiments
For every realization of $\boldsymbol{\omega}$, we made use of the finite element-based solver FEEQS.m [15, 22, 23], developed by H. Heumann and collaborators to obtain a numerical solution of (2). When solving the perturbed problem, the solution corresponding to the baseline values for the currents presented above was used as initial guess for Newton’s iterations.
@@ 4.1 Experiment description
Table 1: Number of spatial grid points $M_{\ell}$ at increasing spatial grid level $\ell=0$ to 5, and number of sparse grid points $P_{q}$ from sparse grid level $q=0$ to 4. In the numerical experiments reported in this paper, we use all spatial levels $\ell=0,\ldots,5$, while the sparse grid level is fixed to $q=1$ due to computational limitations.
@@
For direct non-linear solves at each parameter value, the stopping threshold for the relative residual of Newton’s method is set to $5\times 10^{-11}$. The normalization factor $\|\mathbb{E}[u_{\ell}]\|_{Z}$ on the finest spatial mesh level ($\ell=5$) is approximately $8.5708\times 10^{-1}$. All experiments are conducted using MATLAB R2023a on a System 76 Thelio Major with 256GB RAM and a 64-Core @4.6 GHz AMD Threadripper 3 processor. All experiments are conducted using MATLAB R2023a on a System 76 Thelio Major with 256GB RAM and a 64-Core @4.6 GHz AMD Threadripper 3 processor.
@@ 4.2 Surrogate construction (offline) costs
The cost of constructing the surrogate is then the cost of computing the direct solution on each sparse grid node times the number of sparse grid nodes. For these tests, the latter number is $25$, corresponding to sparse grid level $q=1$. We explored the cost of a single direct solution in [14]. These are summarized in the blue curve shown in the right panel of Figure 2 (This part of the image is reproduced from [14]). They were obtained by solving the systems for 100 random currents and taking the mean CPU time, and they indicate that the costs were of the form $M^{\gamma}$ for $\gamma=1.09$. The resulting total costs are summarized in Table 3 and plotted in the right image of Figure 2, for both single-level and multilevel (in space) surrogates. The plots indicate that the two surrogates have asymptotic complexity of magnitude $\epsilon^{-1}$, which aligns with the theoretical cost $\epsilon^{-\gamma/\alpha}\approx\epsilon^{-1.09}$. The slightly larger cost of the multilevel version is due to the fact that it requires the surrogates for all coarser grids, as observed in the comments that follow (20).
@@
Figure 2: Left: Mean CPU time per realization vs. number of spatial grid points $M_{\ell}$ for both direct and surrogate computations. The computing time was obtained by averaging over 100 realizations. Right: Offline costs of construction of both single spatial level and multi spatial level surrogates.
@@
4.3 Memory considerations
In addition to computational time, it is important to compare the memory requirements of the surrogate construction with those of the direct solver. For the P1 finite element discretization considered here, the stiffness matrices are sparse with a uniformly bounded number of non-zero entries per row under standard mesh regularity assumptions. Consequently, the memory required for a single direct solve at level $\ell$ scales linearly with the number of spatial degrees of freedom $M_{\ell}$. For surrogate construction, additional memory costs arise from storing the direct solutions at each sparse grid node. If $P_{q}$ denotes the number of sparse grid nodes and $L$ the finest spatial level required to meet a given tolerance, then the memory required for the single-level spatial surrogate scales as $\mathcal{O}(P_{q}M_{L})$. In the multilevel case, solutions on all levels $\ell=0,\dots,L$ must be retained, yielding a memory requirement of $\mathcal{O}\!\left(P_{q}\sum_{\ell=0}^{L}M_{\ell}\right)$. Since $M_{\ell}$ grows geometrically with $\ell$, this sum is dominated by $M_{L}$, and thus the multilevel memory cost is also asymptotically $\mathcal{O}(P_{q}M_{L})$.
In the present experiments, the sparse grid level is fixed at $q=1$, so $P_{q}=25$ remains constant as the tolerance $\epsilon$ decreases. Accordingly, the memory usage increases only through the growth of $M_{L}$ as finer spatial grids are required (Table 2). The surrogate therefore introduces a moderate multiplicative memory factor proportional to $P_{q}$, while preserving the same asymptotic scaling with respect to spatial resolution as the direct approach.
@@ 4.4 Sampling (online) costs
The MLMC-FE estimator for (3) involves accumulating surrogate sample corrections $\widehat{Y}_{\ell}^{(i)}$ across $L$ spatial levels. As we discussed in previous work [14], for each sample correction, $\widehat{u}_{\ell-1}^{(i)}$ is constructed on the coarse spatial grid by interpolating $\widehat{u}_{\ell}^{(i)}$ from the fine grid. This interpolation method serves as a more efficient, though slightly less accurate, alternative to the Galerkin projection. However, the use of non-nested geometry-conforming uniform spatial meshes introduces extrapolation errors during interpolation. This issue can be addressed using a sufficiently fine common grid (e.g., $\ell=5$) that encompasses all coarser meshes, in order to minimize or eliminate extrapolation errors caused by interpolation of surrogate corrections from all multi-level spatial coarser grids. Although this does increase accuracy, it does so at a significant extra cost.
Table 4 provides a quantitative summary of the online CPU time for the various solution strategies, including interpolation to a common fine grid just discussed. The discrete solution corresponding to the baseline currents presented at the beginning of Section 4 was used as the initial guess for all non-linear iterations. Since all samples share the same geometry, mesh, and finite element spaces, the associated data structures and assembly routines are initialized once at the beginning of the computation. Thus, the reported solution times reflect primarily the nonlinear solve and linear system solution for each realization. The one-time setup cost is small compared with the cumulative cost of the many solves required for Monte Carlo or multilevel sampling, and is therefore effectively amortized and negligible in the regimes considered here. The entries in the table correspond to the time required to make the square root of the sum of the squares of the discretization and statistical errors (see (22)) less than the tolerances, since the interpolation error is limited by the fixed choice of a sparse grid. The table uses the Monte Carlo method with direct computation as a benchmark and calculates the speedups of the different sampling strategies. It is evident that simply replacing the direct solve with the surrogate (third column) produces significant speedups (on the order of 10 to 50); replacing the full grids with multilevel grids (fourth column) results in somewhat better speedups (up to 200 for small tolerance), and using both surrogates and multilevel grids (fifth column) yields dramatic speedups, often exceeding $10^{4}$. The last two columns of the table present results for mitigating extrapolation errors using a common fine grid for MLMC with both direct solve and surrogate, revealing that the speedups for both approaches are comparable due to the significant interpolation cost. In the following, we discuss some issues regarding accuracy.
Timings without post–processing Timings with post–processing
@@
where the term $25C_{d}$” reflects the overhead of the offline costs for the level-1 collocation sparse grid. This means that the surrogate is less expensive if $k>(21\cdot 25)/20=26.25$. It is clear from Table 5 that typically many more samples are needed in practice. This comparison will also be affected by many factors: the direct solve entails solution of non-linear systems, which will be affected by the number of iterations, and as shown in Table 5, the grid sizes and number of samples will not in general be identical for the different strategies. Nevertheless, this simplified analysis clearly indicates the advantages of the surrogate for a simulation, despite its offline overhead.
@@ 4.5 Properties of geometric parameters
$\mathbf{M}:=\int_{\Omega}\phi_{i}\phi_{j}\,dr\,dz\qquad\text{ and }\qquad\mathbf{S}:=\int_{\Omega}\nabla\phi_{i}\cdot\nabla\phi_{j}\,dr\,dz$
@@ 6 Acknowledgements
Statement of code availability
The code used for the numerical experiments in this article has three components:
  1. 1.
    The free–boundary solver used to obtain the “truth” solutions for the construction of the surrogate, FEEQS.m [22] is a Matlab implementation of the code CEDRES++ [23], kindly provided by Holger Heumann. The authors of this communication are not at liberty of sharing the source. However, the interested reader is referred to the contact information found at [22, 23] for code requests.
  2. 2.
    The code producing the sparse grid surrogates SPINTERP, due to Klimke and Wohlmuth, is available from the Association of Computing Machinery’s Transactions on Mathematical Software as the supplementary material for the article [27].
  3. 3.
    The routines used to interface between SPINTERP and FEEQS.m, the implementation of Algorithm 1, and the post–processing tools used to extract geometric information, perform heat flow regularization, statistical analysis, visualization, etc. are due to the authors of this communication and can be made available upon request.